1. ABCD is rectangle, in which AB = CD = 100 mm and BC = DA = 80 mm. Forces of 100 N each act along AB and CD and forces of 50 N each at along BC and DA. Find the resultant moment of the two couples.?
1
Expert's answer
2021-10-29T02:33:12-0400


Skyli provides a workable method. I prefer a method easier to visualize:

"2 \times 100 \times 40 Nmm + 2 x 50 \times 50 N.mm" "= 13.0 \times newton.meters"


as though the couple were acting through the centroid.